OpenStudy (anonymous):

Write the equation of the line that is parallel to the line 3x - y = -3 and passes through the point (4, -2).

OpenStudy (anonymous):

3x-y=14

OpenStudy (anonymous):

so it would be y = 3x - 14 ?

OpenStudy (anonymous):

yes both are same

OpenStudy (anonymous):

Okay that is what I was thinking, just wanted to double check thanks @vipul92

OpenStudy (anonymous):

0 3x-y=-3 y=3x+3 so the slope of the function is 3 let f(x)=mx+b, m=3, x=4, f(4)=-2, -2=3(4)+b b=-14 so the function is y=3x-14